OpenStudy (anonymous):

Could use the extra credit :) Prove that for a rectangle of fixed perimeter, the square gives the largest area and that given a fixed area, the square gives the smallest perimeter. I proved the first, don't know how to prove the second. It's a corollary. P = 2l + 2w P = 2l+ 2x (given that w=x) P-2x =2l l = (P-2x)/2 A= lw =((P-2x)/2)(x) =(1/2)(Px-2x^2) dA/dt = (1/2)(P-4x) 0 =(1/2)(P-4x) --- Stationary points = max or min 0=P-4x x=P/4 For the area of a fixed P in a rectangle to be largest, each side must be 1/4 the perimeter, therefore a square. Help me prove the 2nd pt of the proof plz

OpenStudy (asnaseer):

in your proof above, you should also show that the 2nd derivative is negative for the value of x that makes the 1st derivative zero. that will prove it is a maximum and not a minimum. the 2nd proof is very similar: A=lw so l=A/w P=2l+2w=2A/w + w so Pw = 2A + w^2 find dP/dw and then see what value of w makes this zero, etc - same principal as you followed above.